# RDMA-Based Algorithms for Sparse Matrix Multiplication on GPUs

### v1

@@ V-C Data Access Primitives
In terms of data access primitives, all direct manipulation of data is performed using put and get RDMA operations. To retrieve a tile of the dense matrix, a process will issue an RDMA get operation to retrieve all or part of a remote tile. Similarly, to write to the tile, a process can issue a remote put operation. In the sparse case, reading from a tile involves issuing remote get operations to retrieve each of the value, row pointer, and column index arrays in a CSR data structure. Due to the complexities involved in modifying a CSR data structure in place, only the process that owns a sparse tile is permitted to modify it. This is done by calling a function replace_tile() to replace the old tile with a new one, followed by a collective function renew_tiles() that will make all sparse tile modifications visible to other processes. As discussed in Section III-A2, when a process needs to send an update to a remote tile, as in the A and B stationary algorithms, we use a distributed queue to send a pointer to the update to the process that owns a particular tile, who will then perform an accumulation at the destination tile. For the update queue, we use BCL’s CheckSumQueue, which enqueues a pointer using a single fetch-and-add operation and an RDMA put, while allowing simultaneous enqueues and dequeues.
@@
As discussed in Section III-A1, the primary primitive for accessing remote tiles of the matrix is get_tile(), which fetches a remote tile of the matrix, and has the same API for both dense and sparse matrices. In the asynchronous version, we return a future object, which will return the local object when the method get() is called. In order to ensure the best possible performance and simplify memory management, we allocate most of each GPU’s memory as shared memory that can be addressed remotely by NVSHMEM, and use our own memory allocator to allocate memory when necessary. This (1) ensures that the local destination buffers in remote get operations are allocated in the shared memory segment, which is a requirement for NVSHMEM transfers over Infiniband, and (2) reduces the overhead of memory allocation during the algorithm, since cudaMalloc tends to be much less efficient than using a custom memory allocator.
V-D MPI SUMMA Implementations
@@ VI Evaluation
On Summit, all codes were compiled with GCC 8.1.1, CUDA 11.2.0, and IBM Spectrum MPI 10.3.1.2. On the single-node DGX-2 system, codes were compiled with GCC 10.2.0, CUDA 11.1.0, and Open MPI 4.0.5. Our RDMA-based implementations use NVSHMEM 2.0.2. All performance experiments use CuSPARSE for local matrix computation. In addition to our own SpMM and SpGEMM implementations, we also benchmark CombBLAS’s GPU SpMM and PETSc’s GPU SpGEMM [16, 25]. We corresponded with CombBLAS’s authors on how to build and execute the gpu branch of CombBLAS on Summit, and with OLCF staff for how to properly build PETSc with GPU support.
The sparse matrices in our experiments are detailed in Table I. Load imbalance is defined as max(nnz($A_{i,j}$))/ave(nnz($A_{i,j}$)) where $A_{i,j}$ is the matrix assigned to a single processor $P(i,j)$. We multiply these matrices against dense matrices with 128 and 512 columns. All matrices use 32-bit floating point values, and for indices, we use 32-bit integers except on "Isolates, Subgraph2" and "Friendster," which require 64-bit indices due to their size.
Figure 3: Single-node runtimes for SpMM, with different numbers of columns $N$ in the dense matrix B.
TABLE I: Matrices used in our experiments. “load imb.” lists the imbalance in number of nonzeros when split amongst 100 processes on a $10\times 10$ 2D process grid.
@@ VI-A SpMM Experiments
Figures 3 and 4 show the results for a variety of different algorithms. The RDMA-based algorithms include a stationary C algorithm (“S-C RDMA”), a stationary A algorithm (“S-A RDMA”), stationary A algorithm with random workstealing (“R WS S-A RDMA”), and locality-aware workstealing with stationary A and C distributions (“LA WS S-C RDMA” and “LA WS S-A RDMA”). Bulk synchronous benchmarks include our own CUDA-aware MPI SUMMA implementation (“BS SUMMA MPI”), and CombBLAS’s GPU SpGEMM kernel (“CombBLAS GPU”). We use matrix sizes representing typical workloads used in a range of real-world applications, such as Graph Neural Networks (GNNs), iterative methods, and graph algorithms[13, 14, 15].
@@
Figure 4: Multi-node runtimes for SpMM, with different numbers of columns $N$ in the dense matrix B.
@@
Our RDMA-based workstealing algorithms are able to achieve performance improvements for some of the less load balanced matrices, such as com-Orkut and friendster, particularly at higher node counts, where the smaller amount of work per node is more likely to lead to some nodes being left with uneven amounts of work. The locality-aware workstealing algorithms tend to perform much better than the random workstealing algorithm, which often incurs a high cost when performing workstealing, since it is not guaranteed that any of the involved matrix tiles will be local. In Table II(a), we can see that the locality-aware workstealing algorithm is able to reduce the amount of time lost due to load imbalance compared to the other RDMA implementations. It should be noted that load imbalance numbers for the RDMA implementations cannot be directly compared to the bulk synchronous MPI implementation, since time lost to load imbalance is tied up in MPI’s broadcast operations, which perform communication, but also enforce synchronization.
VI-B SpGEMM Experiments
To evaluate our RDMA-based SpGEMM algorithm, we perform the computation $C=AA$ for a number of different sparse matrices, representative of a number of different graph algorithms, such as Markov clustering [26], triangle counting [17], and transitive reduction [27]. We again benchmark RDMA-based stationary C and A algorithms, along with a work-stealing version, a SUMMA implementation in MPI, and PETSc’s SpGEMM using CuSPARSE.
Strong scaling performance results are shown in Figure 5. Unlike SpMM, both the single- and multi-node environments are more likely to be compute bound. As a result, all the implementations have similar performance, because they are all performing the same computations. The only exception is PETSc, which is significantly slower, probably because it is not utilizing GPUDirect RDMA. On the large isolates matrix, our bulk synchronous CUDA-aware MPI implementation matches that of the RDMA implementation. On mouse gene, which is somewhat smaller and less load-balanced, we observe good scaling until very high concurrencies, but the RDMA-based implementations achieve higher overall performance as well as somewhat better scaling. On nlpkkt160, which has an uneven distribution of nonzeros, we observe poor performance in the multi-node experiments due to imbalance in both communication and computation. However, the RDMA implementations still perform better than the bulk synchronous implementations. Unlike SpMM, we do not meet our SpGEMM model’s roofline bound on Summit, as shown in Figure 2. This is because the local cuSPARSE operations are not meeting the local roofline bound, evidenced by SpGEMM being compute bound in Table II(b). This likely indicates that there are opportunities for optimization in the local cuSPARSE SpGEMM operations.
@@ VI-B SpGEMM Experiments
TABLE II: Component breakdown for selected matrices.
@@
Figure 5: SpGEMM strong scaling experiments.
VII Related Work and Conclusions
The C++ PGAS library DASH, which uses RDMA through MPI’s one-sided communication API, includes dense matrix and vector data structures as well as an implementation of stationary C dense matrix multiplication [24, 28]. However, DASH does not include any support for sparse matrices or GPUs.
The Combinatorial BLAS library (CombBLAS) [29] is an MPI-based library that provides distributed data structures for dense and sparse matrices, also implementing a number of distributed sparse matrix multiply algorithms. While focusing more on tensor computations, Cyclops Tensor Framework (CTF) [30] also supports several distributed sparse matrix multiplication algorithms.
Both CombBLAS and CTF implement also include so-called communication-avoiding algorithms for sparse matrices, which replicate some of the matrices to reduce communication costs. These ideas have also been applied to SpGEMM [31] and SpMM within the context of graph neural network training [14]. Our work on asynchronous, RDMA-based implementations is orthogonal to these techniques, and could be combined.

### v2

@@ 5.3. Data Access Primitives
In terms of data access primitives, all direct manipulation of data is performed using put and get RDMA operations. To retrieve a tile of the dense matrix, a process will issue an RDMA get operation to retrieve all or part of a remote tile. Similarly, to write to the tile, a process can issue a remote put operation. In the sparse case, reading from a tile involves issuing remote get operations to retrieve each of the value, row pointer, and column index arrays in a CSR data structure. Due to the complexities involved in modifying a CSR data structure in place, only the process that owns a sparse tile is permitted to modify it. This is done by calling a function replace_tile() to replace the old tile with a new one, followed by a collective function renew_tiles() that will make all sparse tile modifications visible to other processes. As discussed in Section 3.1.2, when a process needs to send an update to a remote tile, as in the A and B stationary algorithms, we use a distributed queue to send a pointer to the update to the process that owns a particular tile, who will then perform an accumulation at the destination tile. For the update queue, we use BCL’s CheckSumQueue, which enqueues a pointer using a single fetch-and-add operation and an RDMA put, while allowing simultaneous enqueues and dequeues.
@@
As discussed in Section 3.1.1, the primary primitive for accessing remote tiles of the matrix is get_tile(), which fetches a remote tile of the matrix, and has the same API for both dense and sparse matrices. In the asynchronous version, we return a future object, which will return the local object when the method get() is called. In order to ensure the best possible performance and simplify memory management, we allocate most of each GPU’s memory as shared memory that can be addressed remotely by NVSHMEM, and use our own memory allocator to allocate memory when necessary. This (1) ensures that the local destination buffers in remote get operations are allocated in the shared memory segment, which is a requirement for NVSHMEM transfers over Infiniband, and (2) reduces the overhead of memory allocation during the algorithm, since cudaMalloc tends to be much less efficient than using a custom memory allocator.
5.4. MPI SUMMA Implementations
@@ 6. Evaluation
On Summit, all codes were compiled with GCC 8.1.1, CUDA 11.2.0, and IBM Spectrum MPI 10.3.1.2. On the single-node DGX-2 system, codes were compiled with GCC 10.2.0, CUDA 11.1.0, and Open MPI 4.0.5. Our RDMA-based implementations use NVSHMEM 2.0.2. All performance experiments use CuSPARSE for local matrix computation. In addition to our own SpMM and SpGEMM implementations, we also benchmark CombBLAS’s GPU SpMM and PETSc’s GPU SpGEMM (Buluç and Gilbert, 2011; Balay et al., 2021). We corresponded with CombBLAS’s authors on how to build and execute the gpu branch of CombBLAS on Summit, and with OLCF staff for how to properly build PETSc with GPU support.
The sparse matrices in our experiments are detailed in Table 1. Load imbalance is defined as max(nnz($A_{i,j}$))/ave(nnz($A_{i,j}$)) where $A_{i,j}$ is the matrix assigned to a single processor $P(i,j)$. We multiply these matrices against dense matrices with 128 and 512 columns. All matrices use 32-bit floating point values, and for indices, we use 32-bit integers except on "Isolates, Subgraph2" and "Friendster," which require 64-bit indices due to their size.
Figure 3. Single-node runtimes for SpMM, with different numbers of columns $N$ in the dense matrix B.
Table 1. Matrices used in our experiments. “load imb.” lists the imbalance in number of nonzeros when split amongst 100 processes on a $10\times 10$ 2D process grid.
@@ 6.1. SpMM Experiments
Figures 3 and 4 show the results for a variety of different algorithms. The RDMA-based algorithms include a stationary C algorithm (“S-C RDMA”), a stationary A algorithm (“S-A RDMA”), stationary A algorithm with random workstealing (“R WS S-A RDMA”), and locality-aware workstealing with stationary A and C distributions (“LA WS S-C RDMA” and “LA WS S-A RDMA”). Bulk synchronous benchmarks include our own CUDA-aware MPI SUMMA implementation (“BS SUMMA MPI”), and CombBLAS’s GPU SpGEMM kernel (“CombBLAS GPU”). We use matrix sizes representing typical workloads used in a range of real-world applications, such as Graph Neural Networks (GNNs), iterative methods, and graph algorithms(Huang et al., 2020; Tripathy et al., 2020; Hu et al., 2020).
@@
Figure 4. Multi-node runtimes for SpMM, with different numbers of columns $N$ in the dense matrix B.
@@
Our RDMA-based workstealing algorithms are able to achieve performance improvements for some of the less load balanced matrices, such as com-Orkut and friendster, particularly at higher node counts, where the smaller amount of work per node is more likely to lead to some nodes being left with uneven amounts of work. The locality-aware workstealing algorithms tend to perform much better than the random workstealing algorithm, which often incurs a high cost when performing workstealing, since it is not guaranteed that any of the involved matrix tiles will be local. In Table 2(a), we can see that the locality-aware workstealing algorithm is able to reduce the amount of time lost due to load imbalance compared to the other RDMA implementations. It should be noted that load imbalance numbers for the RDMA implementations cannot be directly compared to the bulk synchronous MPI implementation, since time lost to load imbalance is tied up in MPI’s broadcast operations, which perform communication, but also enforce synchronization.
6.2. SpGEMM Experiments
To evaluate our RDMA-based SpGEMM algorithm, we perform the computation $C=AA$ for a number of different sparse matrices, representative of a number of different graph algorithms, such as Markov clustering (van Dongen, 2000), triangle counting (Azad et al., 2015), and transitive reduction (Guidi et al., 2021). We again benchmark RDMA-based stationary C and A algorithms, along with a work-stealing version, a SUMMA implementation in MPI, and PETSc’s SpGEMM using CuSPARSE.
Strong scaling performance results are shown in Figure 5. Unlike SpMM, both the single- and multi-node environments are more likely to be compute bound. As a result, all the implementations have similar performance, because they are all performing the same computations. The only exception is PETSc, which is significantly slower, probably because it is not utilizing GPUDirect RDMA. On the large isolates matrix, our bulk synchronous CUDA-aware MPI implementation matches that of the RDMA implementation. On mouse gene, which is somewhat smaller and less load-balanced, we observe good scaling until very high concurrencies, but the RDMA-based implementations achieve higher overall performance as well as somewhat better scaling. On nlpkkt160, which has an uneven distribution of nonzeros, we observe poor performance in the multi-node experiments due to imbalance in both communication and computation. However, the RDMA implementations still perform better than the bulk synchronous implementations. Unlike SpMM, we do not meet our SpGEMM model’s roofline bound on Summit, as shown in Figure 2. This is because the local cuSPARSE operations are not meeting the local roofline bound, evidenced by SpGEMM being compute bound in Table 2(b). This likely indicates that there are opportunities for optimization in the local cuSPARSE SpGEMM operations.
@@ 6.2. SpGEMM Experiments
Table 2. Component breakdown for selected matrices.
@@
Figure 5. SpGEMM strong scaling experiments.
7. Related Work and Conclusions
The C++ PGAS library DASH, which uses RDMA through MPI’s one-sided communication API, includes dense matrix and vector data structures as well as an implementation of stationary C dense matrix multiplication (Fürlinger et al., 2016; Mössbauer et al., 2018). However, DASH does not include any support for sparse matrices or GPUs.
The Combinatorial BLAS library (CombBLAS) (Azad et al., 2021) is an MPI-based library that provides distributed data structures for dense and sparse matrices, also implementing a number of distributed sparse matrix multiply algorithms. While focusing more on tensor computations, Cyclops Tensor Framework (CTF) (Solomonik et al., 2013) also supports several distributed sparse matrix multiplication algorithms.
Both CombBLAS and CTF implement also include so-called communication-avoiding algorithms for sparse matrices, which replicate some of the matrices to reduce communication costs. These ideas have also been applied to SpGEMM (Azad et al., 2016) and SpMM within the context of graph neural network training (Tripathy et al., 2020). Our work on asynchronous, RDMA-based implementations is orthogonal to these techniques, and could be combined.
@@ 7. Related Work and Conclusions
Acknowledgements.
This work was supported in part by the National Science Foundation Graduate Research Fellowship Program under Grant No. GE 1752814 as well as under National Science Foundation Award No. CCF-1823034. It was also funded in part by the Advanced Scientific Computing Research (ASCR) program within the Office of Science of the DOE under contract number DE-AC02-05CH11231. This work used resources of the National Energy Research Scientific Computing Center (NERSC) at Lawrence Berkeley National Laboratory supported by the Office of Science of the DOE under Contract No. DE-AC02- 05CH11231 as well as Extreme Science and Engineering Discovery Environment (XSEDE) Bridges-2 at the Pittsburgh Supercomputing Center through allocation ASC180051.